Given cos(x)= (-1/10) with 90 degrees I solved this...
cos^2(theta/2)= (1/2)(1+cos(theta))
=((1/2) ((10/10)-(1/10))
=(1/2)(9/10)
=9/20
cos (theta/2)=sqroot (9/20)
= 3/ (sqroot20)
=(3(sqroot20))/20
=(3*2sqroot5)/20
=(3sqroot5)/10
I marked it as NEGATIVE because it is in the second quadrant but it is marked wrong. However, when I submit positive (3sqroot5)/10 it marks it right. Why does it remain positive even though it is in the second quadrant?


Answer by stanbon(48564)

Given cos(x)= (-1/10) with 90 degrees I solved this...
cos^2(theta/2)= (1/2)(1+cos(theta))
=((1/2) ((10/10)-(1/10))
=(1/2)(9/10)
=9/20
cos (theta/2)=sqroot (9/20)
= 3/ (sqroot20)
=(3(sqroot20))/20
=(3*2sqroot5)/20
=(3sqroot5)/10
I marked it as NEGATIVE because it is in the second quadrant but it is marked wrong. However, when I submit positive (3sqroot5)/10 it marks it right. Why does it remain positive even though it is in the second quadrant?
-------
It accepts the positive answer because you are calculating the
cos of (theta/2). Although theta is in the 2nd quadrant, theta/2
is in the 1st quadrant.
================================
Cheers,
Stan H.